OpenStudy (anonymous):

Find f(g(x)) when f(x)= -2x+7 and g(x)= x+2

OpenStudy (anonymous):

In f(x), for every x, plug in g(x) So: f(x)= -2x + 7 g(x)= x + 2 f(g(x))= -2(x+2)+7 Now evaluate that

OpenStudy (anonymous):

So -2x+11 is the answer?

OpenStudy (anonymous):

I think you just simplified wrong: -4 +7

OpenStudy (anonymous):

-2x +3 ?

OpenStudy (anonymous):

Yep!